### v2

# EVM-QuestBench: An Execution-Grounded Benchmark for Natural-Language Transaction Code Generation

Abstract
Large language models are increasingly applied to various development scenarios. However, in on-chain transaction scenarios, even a minor error can cause irreversible loss for users. Existing evaluations often overlook execution accuracy and safety. We introduce EVM-QuestBench, an execution-grounded benchmark for natural-language transaction-script generation on EVM-compatible chains. The benchmark employs dynamic evaluation: instructions are sampled from template pools, numeric parameters are drawn from predefined intervals, and validators verify outcomes against these instantiated values. EVM-QuestBench contains 107 tasks (62 atomic, 45 composite). Its modular architecture enables rapid task development. The runner executes scripts on a forked EVM chain with snapshot isolation; composite tasks apply step-efficiency decay. We evaluate 20 models with 5 independent rounds each and find large performance gaps, with split scores revealing persistent asymmetry between single-action precision and multi-step workflow completion. Code: https://anonymous.4open.science/r/bsc_quest_bench-A9CF/.
1 Introduction
Large language models (LLMs) are increasingly being used to control software and tools (openai2023gpt4; anil2023palm2). Leveraging LLMs for code generation and blockchain transactions is becoming commonplace, but this introduces significant financial risks. Even a small error, such as an incorrect address, unit, or deadline, can result in irreversible losses.
Benchmarks for code understanding and generation need to cover a broad range of programming tasks and reasoning capabilities (chen2021evaluating; lu2021codexglue). Many evaluations still rely on lexical overlap metrics such as BLEU or CodeBLEU (papineni2002bleu; ren2020codebleu). These metrics can reward outputs that appear similar to references but fail to run or fail to meet functional constraints. Benchmarks such as SWE-bench (jimenez2024swebench) focus on real-world software engineering tasks like bug fixing in Python repositories. Due to the difference in test domains, such benchmarks cannot directly reflect a model’s ability to execute transactions in Web2/Web3 environments. Blockchain-specific benchmarks such as Solana Bench (solana2025bench) explore the boundaries of LLM capabilities in Web3 but struggle to provide feedback on the accuracy of natural language understanding and the safety of transaction execution.
Transaction script generation presents a distinct set of failure modes. LLMs must first interpret diverse natural language instructions. The generated code must correctly construct calldata, account for chain-specific units and token decimals, adhere to protocol constraints, and manage dependencies across multiple transaction steps. Even minor deviations can lead to transaction reverts, partial execution, or incorrect state transitions. These characteristics make blockchain automation an ideal domain for execution-based evaluation.
We study natural language transaction script generation on EVM-compatible chains. The target output is a client-side TypeScript module that composes calls to deployed contracts. The module returns unsigned transaction payloads. A standardized runner signs and executes the requests on a forked chain. Validators check receipts and post-state constraints. This design measures end-to-end correctness under execution.
We introduce EVM-QuestBench, an execution-grounded benchmark with two splits. Atomic tasks test single-action precision. Composite tasks test multi-step workflows that require planning, prerequisite handling, and parameter propagation. Composite scoring incorporates a step efficiency factor that penalizes unnecessary steps. EVM-QuestBench contains 107 tasks with a maximum total score of 10,700. This design significantly simplifies benchmark development and maintenance. Creating an atomic task only requires defining the problem and developing a validator. Creating a new composite task only requires updating a JSON file.
We evaluate 20 models under a unified protocol with 5 independent rounds per model. Results show substantial variance. Split scores reveal a persistent capability asymmetry between single-action precision and workflow completion. Several models achieve strong Composite performance with weaker Atomic scores. Several models fail on Composite workflows despite non-trivial Atomic performance. These patterns motivate split reporting for diagnosis.
Our contributions are as follows:
  • We release EVM-QuestBench, a benchmark for natural language to transaction script generation on EVM-compatible chains, with Atomic and Composite splits.
  • We introduce an atomic/composite benchmark paradigm that significantly reduces development costs, especially when using LLMs to assist development.
  • We provide an execution protocol with snapshot isolation, a fixed runner interface, and validator-based scoring over receipts and post-state constraints.
  • We report results on 20 models with 5-round statistical analysis, including confidence intervals and rank-order consistency, that separates single-action precision from multi-step workflow completion.
2 Related Work
Execution-based evaluation for code generation.
A major limitation of reference-based code evaluation is that surface overlap can reward outputs that resemble references but fail functionally. Execution-based benchmarks address this by running generated programs against tests, including HumanEval and MBPP for function synthesis (chen2021evaluating; austin2021program), and extensions to library usage and multilingual settings (wang2023odex; khan2024xcodeeval; yan2024codescope). However, most prior benchmarks focus on stateless, sandboxed functions and do not model shared external state or irreversible actions.
Real-world software engineering and multi-step interaction.
SWE-bench evaluates issue resolution by applying patches to real repositories and validating with test suites (jimenez2024swebench), emphasizing repository-scale context and cross-file edits. Separately, agent-style benchmarks study multi-step tool use and action sequencing in interactive environments (qin2023toollm; liu2023agentbench). These lines are complementary, but they typically do not target blockchain-specific constraints such as transaction construction, revert conditions, strict unit handling, or protocol-imposed prerequisites.
Blockchain-specific evaluation and our positioning.
Blockchain interaction introduces shared mutable state, protocol prerequisites, and revert risk that are absent in standard code tasks. Prior blockchain benchmarks (e.g., Solana-focused transaction evaluations) (solana2025bench) do not explicitly disentangle single-transaction precision from multi-transaction workflow completion under a unified execution and validation interface. EVM-QuestBench fills this gap by evaluating natural-language-to-transaction-script generation on EVM-compatible chains with atomic/composite splits. Tasks are specified declaratively as JSON with reusable validator components, enabling new tasks to be added with minimal engineering effort while preserving comparable execution-grounded evaluation.
3 Benchmark Tasks
3.1 Benchmark Overview
EVM-QuestBench evaluates whether a model can convert a natural language goal into an executable outcome on EVM-compatible chains. Each task specifies a target end state. For evaluation, the runner samples a template from a pre-built pool and instantiates numeric parameters within predefined ranges. The model outputs a TypeScript module that constructs transaction request objects. The runner executes the plan on a forked chain, and validators score post-execution constraints.
Figure 1 illustrates the layered architecture. The design is modular: adding a new atomic task requires only a JSON specification and a validator; adding a composite task requires only a JSON file. At evaluation time, the system flows through dynamic instantiation (template and parameter sampling), LLM interaction (code generation), code execution (sandboxed TypeScript runtime), transaction execution (signing and broadcasting on the fork), and validation (weighted scoring against dynamically sampled parameters).
Figure 1: EVM-QuestBench evaluation architecture. Tasks are defined as JSON specs, dynamically instantiated into concrete instances, executed through a fixed TypeScript interface, and scored by validator-driven post-state checks. Composite tasks additionally apply step-efficiency decay.
The benchmark targets transaction script generation rather than contract synthesis. Failures commonly arise from incorrect calldata construction, unit conversion, or missing protocol prerequisites.
3.2 Task Specification and Data Format
Each task is stored as a JSON specification containing an identifier, metadata, natural language templates, parameter definitions, and a validation configuration. Atomic and composite tasks share a common surface structure (templates, parameters, and validation). Composite tasks additionally include an explicit workflow structure and a scoring strategy, enabling evaluation of multi transaction dependencies.
Atomic task schema.
An atomic task defines a single on chain action and its expected effects. The specification includes: (i) natural language templates that render user instructions, (ii) typed parameters with default ranges, and (iii) a validator class with task specific arguments (e.g., token address, recipient address, amount, decimals). At evaluation time, the runner instantiates the validator and computes a weighted score from post execution checks.
Composite task schema.
A composite task defines a multi transaction workflow and an end state condition. Each composite specification includes a composite_structure field that names a workflow pattern, sets optimal_steps, and enumerates step level atomic operations. Composite validation checks whether the final on chain condition holds, then applies a step efficiency decay described in Section 4. This design prioritizes end to end completion over intermediate surface matching.
Running example.
A typical instance provides an instruction such as “Swap 0.1 ETH to USDT” together with an execution context (RPC endpoint, agent address, and a contract address map). The model returns a module whose executeSkill function emits a transaction request that calls the router with a concrete to address and data calldata. The validator then checks transaction success and verifies the expected balance change within the task tolerance (Section 3).
3.3 Benchmark Composition
EVM-QuestBench contains 107 tasks (62 atomic, 45 composite) with a maximum total score of 10,700. Figure 2 shows the split sizes and Figure 3 shows the composite workflow complexity.
Figure 2: Task split in EVM-QuestBench: 62 atomic tasks and 45 composite tasks (107 total).
Figure 3: Distribution of optimal steps ($K_{\text{opt}}$) for composite tasks. Most workflows require 3 steps (53.3%), with a mean of $3.$
Atomic tasks require one on-chain action. The task bank is organized into three categories: basic_transactions (40), defi_operations (19), and advanced_features (3). Basic transactions cover wallet and token operations (queries, native transfers, ERC20 transfers, approvals). DeFi operations cover swaps, liquidity, and staking. Advanced features cover edge cases (fallback handling, delegatecall, flashloans). Atomic tasks stress parameter correctness, unit handling, and precise state change targets.
[…]
3.4 Difficulty and Coverage
Each task is annotated with a difficulty label from {easy, medium, hard}. Atomic tasks are dominated by medium (56.5%), while composite tasks skew harder (hard: 24.4%), reflecting additional constraints from prerequisite handling and step ordering. Atomic tasks span subcategories including ERC20 operations, native transfers, NFT operations, swaps, staking, and queries. Composite tasks diversify by workflow motifs: batch operations, swap-centric workflows, liquidity workflows, staking workflows, and query-and-verify patterns.
3.5 Instruction Templates and Parameterization
Each task provides multiple natural language templates that render a user goal as an instruction, together with a parameter schema that binds slots in the template (e.g., token symbols, contract keys, recipient addresses, and amounts). These templates are generated by an LLM during benchmark development to ensure diverse and realistic phrasing. Atomic tasks include 3355 templates (mean $3.97$, median $4$), while composite tasks include 2244 templates (mean $2.82$, median $3$).
At evaluation time, the runner randomly selects one template from the candidate pool and samples numeric parameter values uniformly within predefined ranges specified in the task configuration. For both atomic and composite tasks, all numeric values (e.g., transfer amounts, token quantities, percentages) are dynamically generated within configured intervals rather than fixed. The runner injects these sampled values into the selected template and passes the resolved parameters to validators.
Benefits of dynamic parameterization.
Dynamic numeric sampling provides several evaluation advantages over fixed test cases. First, it prevents trivial memorization: models cannot succeed by caching specific amount patterns from training data. Second, it tests numeric reasoning robustness: models must correctly parse arbitrary values from natural language (e.g., “transfer 0.37 tokens” vs. “transfer 100 tokens”) and propagate them through unit conversions and calldata encoding. Third, it exposes edge-case failures: boundary values and unusual magnitudes reveal precision errors, overflow handling issues, and decimal truncation bugs that fixed test cases might miss. Fourth, it enables statistical significance: multiple runs with different parameter samples yield confidence intervals rather than single-point estimates. This setup ensures that models must generalize across different numeric inputs rather than pattern-match against known examples. Correctness depends on consistent use of task-provided context, including the contract address map, token decimals, and protocol constraints.
3.6 Validators and Post Execution Constraints
EVM-QuestBench uses validator-based scoring rather than reference code matching. This design enables evaluation of functional correctness under dynamic parameterization: validators receive the same randomly sampled parameters that were injected into the natural language instruction, and verify that the model’s output produces the expected on-chain effects for those specific values.
Validator architecture.
Each atomic task type has a dedicated validator class (e.g., ERC20TransferValidator, SwapExactETHForTokensValidator). At instantiation, validators receive the dynamically sampled parameters from the runner, including target addresses, token amounts, and decimal configurations. Composite tasks use a CompositeValidator that loads and delegates to the appropriate atomic validator based on the key operation in the workflow.
Validation against dynamic parameters.
The validator converts human-readable parameters to chain-native units (e.g., $0.1$ tokens $\rightarrow$ $0.1\times 10^{18}$ wei for an 18-decimal token) and compares against actual on-chain state changes. For a transfer task, the validator checks: (i) the sender’s balance decreased by the expected amount, and (ii) the receiver’s balance increased by the same amount. This approach ensures that correctness is measured against the specific numeric values presented in the instruction, not against a fixed reference output.
Weighted scoring components.
Validators implement protocol and task-specific checks over receipts and post-state. A typical atomic validator (e.g., ERC20 transfer) computes a weighted score from four check categories: transaction success (30 points), contract address correctness (20 points), function signature correctness (20 points), and state change verification (30 points). Each check is binary pass/fail with its assigned weight. The final score is the sum of passed check weights.
Tolerance handling.
Many tasks allow small tolerances to account for AMM slippage or rounding effects. Transfer validators typically allow a $1$ to 22% relative tolerance on amounts. Swap and liquidity operations use a slippage-style tolerance (commonly 55%). Approval tasks require an exact allowance match. These tolerances are applied during the state change verification step, ensuring that minor deviations from expected values (due to protocol mechanics) do not penalize otherwise correct solutions.
Composite validation.
Composite validators compute a base success score from the final end-state condition, then apply step efficiency decay as described in Section 4. The validator loads the composite task definition, identifies the key operation, and reuses the corresponding atomic validator to score the final transaction. This modular design eliminates code duplication and ensures scoring consistency between atomic and composite evaluations of the same operation type.
3.7 Score Reporting
We report three aggregate scores per model: Atomic score (sum over 62 tasks), Composite score (sum over 45 tasks), and Total score (sum over 107 tasks). We also report per task averages by normalizing each split sum by its task count. For summary statistics, we use a pass threshold of 60 points per task.
4 Evaluation Setup
We treat correctness as an end-to-end property: a submission must construct valid transaction request objects, execute successfully in a forked environment, and satisfy post-state constraints checked by validators.
4.1 Task Formalization
Atomic tasks.
Let $\mathcal{A}=\{a_{1},a_{2},\ldots,a_{n}\}$ denote the set of atomic tasks. Each atomic task $a_{i}$ is defined as a tuple:
$a_{i}=(T_{i},P_{i},V_{i},S_{i}^{\max})$
where $T_{i}$ is the set of natural language templates (one is randomly selected at test time), $P_{i}$ is the parameter space (numeric values are sampled uniformly within predefined intervals), $V_{i}$ is the validator that checks post-execution state changes, and $S_{i}^{\max}=100$ is the maximum score.
Composite tasks.
A composite task $C$ is constructed from an ordered sequence of atomic operations:
$C=\bigl(a_{i_{1}}\rightarrow a_{i_{2}}\rightarrow\cdots\rightarrow a_{i_{m}},\ K^{\mathrm{opt}},\ V_{C}\bigr)$
where $a_{i_{j}}\in\mathcal{A}$ is the atomic operation at step $j$, $K^{\mathrm{opt}}$ is the optimal number of steps specified by the task definition, and $V_{C}$ is the composite validator that checks the final state condition. The arrow notation $\rightarrow$ indicates execution order and potential data dependencies (e.g., an approval must precede a swap that requires that allowance).
4.2 Execution Environment
Evaluation runs on an Anvil fork of BSC mainnet (chain ID 56), an EVM-compatible chain. We use BSC as the instantiation; the benchmark design generalizes to other EVM chains. The runner specifies the upstream RPC endpoint. The fork block height is not pinned; each evaluation uses the latest block at fork time. Within an evaluation run, we enforce comparability across tasks via snapshot isolation (Section 4.3).
Each run creates a fresh test account. The private key is generated at runtime and is never exposed to the model. Transaction signing happens in the evaluation process, separating generation from authorization and preventing the model from directly controlling keys.
Before tasks start, the account is funded with 100 BNB. The environment provisions task assets, including ERC20 tokens, LP tokens, and NFT holdings. The runner also deploys a small set of auxiliary contracts at runtime for testing.
4.3 Isolation
After environment initialization, we create a snapshot of the forked chain state. Before executing each task, the runner restores the snapshot, yielding a consistent initial state per task. This prevents cross-task interference and makes task scores comparable under identical starting conditions.
4.4 Inference and Subtask Planning
Unless otherwise specified, we use single-shot generation for atomic tasks: each model is called once per task instance to produce a complete TypeScript module. We set the decoding temperature to $0.7$.
For composite tasks, we employ a multi-turn interaction protocol with explicit subtask planning. The LLM first enters a planning phase, where it analyzes the natural language instruction and decomposes the task into an ordered sequence of subtasks. Formally, the model outputs a plan:
$\Pi=\langle\pi_{1},\pi_{2},\ldots,\pi_{k}\rangle$
where each subtask $\pi_{j}=(t_{j},c_{j},p_{j})$ specifies an action type $t_{j}$ (e.g., approve, swap, stake, query), target contract $c_{j}$, and parameters $p_{j}$. The planning objective is to minimize the total number of steps while satisfying task constraints:
$\min|\Pi|\quad\text{subject to}\quad\Pi\models C$
where $\Pi\models C$ denotes that executing $\Pi$ achieves the goal state specified by composite task $C$.
After planning, the LLM enters the execution phase, iterating through the planned subtasks. For each subtask, the model generates a TypeScript code block that returns a transaction request object. The runner executes the transaction on the forked chain and returns the result (success/failure, receipt, state changes) to the LLM. The model can also perform query actions to inspect on-chain state (balances, allowances) before executing transactions. This closed-loop interaction continues until the model signals task completion or exhausts the round budget.
4.5 Scoring
Each task has a maximum score of $S^{\max}=100$.
Atomic scoring.
Atomic tasks use task-specific validators. Each validator computes a weighted score from a set of post-execution checks $\mathcal{C}$:
$S_{\mathrm{atomic}}=\sum_{c\in\mathcal{C}}w_{c}\cdot\mathbf{1}\bigl[f_{c}(s_{\mathrm{pre}},s_{\mathrm{post}})\bigr]$
where $w_{c}$ is the weight assigned to check $c$, $f_{c}$ is the check function comparing pre-execution state $s_{\mathrm{pre}}$ and post-execution state $s_{\mathrm{post}}$, and $\mathbf{1}[\cdot]$ is the indicator function. The weights satisfy $\sum_{c\in\mathcal{C}}w_{c}=100$. Typical checks include transaction success (30 points), target address correctness (20 points), function signature correctness (20 points), and state-change verification through balance or allowance deltas (30 points). Where required by on-chain mechanics, validators apply tolerances to account for rounding or protocol-side price impact.
Composite scoring.
Composite tasks use outcome-based scoring with step-efficiency decay. Let $K^{\mathrm{act}}$ denote the actual number of execution rounds (including failed attempts and retries up to the retry budget). The final score is:
$S=S_{\mathrm{base}}\cdot\min\!\left(1,\ \frac{K^{\mathrm{opt}}}{K^{\mathrm{act}}}\right)$
where $S_{\mathrm{base}}\in\{0,100\}$ depends on whether the final state condition holds. This formulation rewards efficient execution: if $K^{\mathrm{act}}\leq K^{\mathrm{opt}}$, the model receives full credit; if $K^{\mathrm{act}}>K^{\mathrm{opt}}$, the score decays proportionally to the ratio of optimal to actual steps.
Aggregate scoring.
Let $\mathcal{A}_{\mathrm{bench}}$ denote the set of 62 atomic tasks and $\mathcal{C}_{\mathrm{bench}}$ denote the set of 45 composite tasks. We report:
$S_{\mathrm{Atomic}}$ $=\sum_{a\in\mathcal{A}_{\mathrm{bench}}}S_{a},$
$S_{\mathrm{Composite}}$ $=\sum_{c\in\mathcal{C}_{\mathrm{bench}}}S_{c},$
$S_{\mathrm{Total}}$ $=S_{\mathrm{Atomic}}+S_{\mathrm{Composite}}$
The maximum possible scores are $S_{\mathrm{Atomic}}^{\max}=6{,}200$, $S_{\mathrm{Composite}}^{\max}=4{,}500$, and $S_{\mathrm{Total}}^{\max}=10{,}700$.
4.6 Artifacts
We release the benchmark codebase, task definitions, validators, and the runner. We release aggregate model scores. We do not release full per-task outputs, JSON logs, or execution traces at this time.
5 Results and Analysis
5.1 Leaderboard
We evaluate 20 models on EVM-QuestBench with 5 independent evaluation rounds per model (totaling $5\times 107\times 20=10{,}700$ task executions). Each round uses freshly sampled numeric parameters, independent LLM calls (temperature =0.7=0.7), and clean snapshot-isolated environments. Table 1 reports the mean Atomic, Composite, and Total scores across 5 rounds, along with standard deviation (SD), 95% confidence interval (CI, computed as mean $\pm\,t_{0.025,4}\times\text{SE}$ with $t=2.776$ for df =4=4), and coefficient of variation (CV%).
The top three models exceed 7,7007{,}700 mean total points. Claude-Sonnet-4.5 achieves the highest mean total (8,2368{,}236) with low variance (CV $2.1\%$). 15 out of 20 models exhibit CV% below $9\%$, demonstrating that evaluation scores are robust to the combined variance from dynamic parameter sampling, temperature-based generation, and RPC execution.
# Model Atom Comp Total SD CV% 95% CI
1 Claude-Sonnet-4.5 4278 3958 8236 174 2.1 [8020, 8451]
2 Gemini-3-Pro 4302 3561 7863 267 3.4 [7531, 8195]
3 GPT-5 4128 3646 7774 407 5.2 [7269, 8279]
4 GPT-5.1 3715 3617 7332 144 2.0 [7154, 7510]
5 Kimi-K2-Thinking 3605 3633 7238 207 2.9 [6982, 7495]
6 Gemini-2.5-Flash 3265 3768 7033 127 1.8 [6876, 7191]
7 GPT-OSS-120B 3458 3468 6926 436 6.3 [6384, 7467]
8 DeepSeek-V3.2 3071 3638 6709 384 5.7 [6232, 7186]
9 Grok-4-Fast 2980 3207 6187 224 3.6 [5909, 6465]
10 Qwen3-235B 3609 2116 5724 630 11.0 [4943, 6506]
11 GPT-5.1-Mini 2510 3190 5700 510 8.9 [5067, 6333]
12 Qwen3-Max 2775 2883 5658 229 4.1 [5374, 5943]
13 Claude-Haiku-4.5 3221 2367 5588 316 5.7 [5196, 5981]
14 Mimo-V2-Flash 3189 1723 4912 218 4.4 [4641, 5183]
15 GLM-4.6 2881 1971 4852 295 6.1 [4486, 5218]
16 Qwen3-30B 1554 2011 3565 398 11.2 [3072, 4059]
17 Devstral-2512 2892 21 2914 303 10.4 [2537, 3290]
18 Qwen3-Coder 1801 160 1961 291 14.8 [1600, 2322]
19 Qwen3-Coder-Flash 223 38 260 212 81.4 [3-3, 524]
20 Qwen3-Coder-30B 140 5 145 61 42.0 [69, 221]
Table 1: Leaderboard (avg@5). All scores are means over 5 independent rounds. Atom = Atomic (62 tasks, max 6,200), Comp = Composite (45 tasks, max 4,500), Total = 107 tasks (max 10,700). Spearman’s rank correlation between any two rounds averages $\rho=0.960$ (all $p<0.001$, range: 0.940–0.983).
5.2 Atomic versus Composite Capability Separation
Figure 4 visualizes the Atomic–Composite plane (a full ranking bar chart is in Appendix E, Figure 6). Models spread widely, indicating that single-step correctness and multi-step completion are only partially coupled.
We observe two characteristic asymmetries. Workflow-oriented models (DeepSeek-V3.2, Gemini-2.5-Flash, GPT-5.1-Mini) achieve high Composite despite weaker Atomic, suggesting stronger sequencing and end-state targeting. Precision-oriented models (Claude-Haiku-4.5, Devstral-2512) achieve high Atomic but lag on Composite, consistent with weaknesses in multi-step dependency tracking. Several code-specialized models (Qwen3-Coder-30B, Devstral-2512, Qwen3-Coder) score near zero on Composite due to repeated interface failures in multi-step workflows.
Figure 4: Atomic score versus Composite score (avg@5). Each point is a model.
5.3 Composite Execution Quality
Table 2 reports fine-grained composite metrics averaged over 5 rounds. We analyze step efficiency ($\bar{K}_{\text{act}}$, Eff%) as primary dimensions.
Model Score Pass $\bar{K}$ Eff%
Claude-Sonnet-4.5 3958 41.0 3.8 88.0
Gemini-2.5-Flash 3768 38.6 4.1 83.7
GPT-5 3646 38.0 3.6 83.8
DeepSeek-V3.2 3638 38.2 3.5 84.5
Kimi-K2-Thinking 3633 33.6 3.6 84.2
GPT-5.1 3617 38.0 4.1 80.4
Gemini-3-Pro 3561 37.6 4.3 79.7
GPT-OSS-120B 3468 35.2 3.9 82.6
Grok-4-Fast 3207 32.0 4.4 76.4
GPT-5.1-Mini 3190 31.0 4.7 73.6
Qwen3-Max 2883 25.8 4.9 71.5
Claude-Haiku-4.5 2367 13.0 5.9 60.7
Qwen3-235B 2116 11.0 5.8 62.6
Qwen3-30B 2011 8.4 6.1 58.8
GLM-4.6 1971 10.2 6.0 58.3
Mimo-V2-Flash 1723 9.0 6.2 56.9
Qwen3-Coder 160 0.4 6.7 53.3
Qwen3-Coder-Flash 38 0.2 6.8 51.1
Devstral-2512 21 0.0 6.8 50.8
Qwen3-Coder-30B 5 0.0 6.8 50.6
Table 2: Composite execution quality (avg@5, 45 tasks). Score and Pass are means over 5 rounds. $\bar{K}$ = mean $K_{\text{act}}$, Eff = step efficiency.
Step efficiency reflects planning quality.
Claude-Sonnet-4.5 leads with 88.0% efficiency and the highest composite score (3,958 avg), completing an average of 41 out of 45 tasks. DeepSeek-V3.2 achieves the lowest $\bar{K}_{\text{act}}$ (3.5) with 84.5% efficiency, indicating consistent per-task optimization. Lower-tier models average $\bar{K}_{\text{act}}>5.9$, which compounds both failure risk and score decay.
Pass count separates tiers.
Top models pass 33–41 of 45 composite tasks on average across 5 rounds. Claude-Haiku-4.5 passes only 13 tasks despite reasonable atomic performance, confirming that multi-step workflow completion requires capabilities beyond single-action precision. Four models (Devstral-2512, Qwen3-Coder-30B, Qwen3-Coder-Flash, Qwen3-Coder) pass fewer than 1 task on average, reflecting systematic interface or planning failures in multi-step workflows.
Interface compliance.
All 20 models produce syntactically valid TypeScript modules that the runner can execute on atomic tasks. However, three code-specialized models (Qwen3-Coder-30B, Devstral-2512, Qwen3-Coder) score near zero on composite tasks due to repeated schema errors such as missing ethers imports or incorrect module structure in multi-step contexts. We retain these models in the leaderboard to document the full spectrum of interface-level and capability-level failures.
5.4 Where Workflows Fail
Figure 5 reports pass rates by workflow pattern. The hardest patterns are multi stage DeFi workflows combining liquidity and staking, as well as batch approval patterns. These concentrate three failure surfaces: (i) prerequisite correctness, (ii) cross step parameter consistency, and (iii) execution robustness under multiple transactions. Additional analysis is in Appendix E.
Figure 5: Composite workflow difficulty by pattern. Pass is defined as score $\geq 60$.
6 Conclusion
We presented EVM-QuestBench, an execution-grounded benchmark for natural language transaction script generation on EVM-compatible chains, instantiated on BNB Smart Chain (chain ID 56). The benchmark contains 107 tasks with Atomic and Composite splits. Evaluation executes model generated TypeScript scripts in a forked environment with snapshot isolation. Validators score outcomes after execution. Composite tasks apply a step efficiency factor.
We evaluated 20 models under a unified runner with 5 independent evaluation rounds per model. Performance varies substantially across models. Split scores expose a persistent gap between single transaction precision and workflow completion. Several models achieve strong Composite scores with weaker Atomic scores. Several models fail on Composite tasks despite non zero Atomic performance.
EVM-QuestBench offers a standardized protocol for studying execution-grounded behavior in on-chain automation. Future work can expand task coverage, extend the benchmark to additional EVM chains, and strengthen analysis using per task traces. In fact, we have already leveraged the same architecture to develop and establish a benchmark on Solana, demonstrating the portability of our atomic/composite paradigm across heterogeneous blockchain ecosystems. We also plan to incorporate richer security checks for transaction intent and side effects. Additionally, we intend to evaluate LLMs’ success rates in generating atomic and composite task definitions, which would further assess their understanding of on-chain transactions and smart contract semantics.
7 Limitations
Execution stability.
Because EVM-QuestBench is execution-grounded, scores inherit instability from RPC connectivity, fork performance, and provider availability. Snapshot isolation and composite retries reduce, but do not eliminate, this effect.
Reproducibility across forks.
By default, the fork block height is not pinned, and cross-run results can drift with mainnet state, especially for price-sensitive swaps and liquidity operations. Our framework supports pinning the fork at a specific block height via a configuration file; we recommend this setting for standardized comparative benchmarking. Recording the fork block number in evaluation metadata further aids reproducibility.
Scoring and interface effects.
Composite scoring applies a step efficiency decay factor $\min(1,K_{\mathrm{opt}}/K_{\mathrm{act}})$, where $K_{\mathrm{act}}$ counts execution rounds that may include retries; this can conflate planning inefficiency with execution instability. Validators also use tolerances for some tasks, which may admit occasional false positives or over penalize benign deviations. Finally, providers differ in default output budgets and formatting behavior, which can affect schema validity; prior work has noted that automatic metrics and LLM as a judge can introduce additional biases (zheng2023judging; liu2023geval).
Number of evaluation rounds.
We conduct 5 independent evaluation rounds per model and report mean scores with confidence intervals (Table 1). While 5 rounds provide reasonable statistical power for tier-level separation, some adjacent-rank differences remain statistically indistinguishable (overlapping CIs). Additional rounds would narrow confidence intervals and enable finer-grained ranking distinctions.
Appendix A Reproducibility
This appendix summarizes the setup required to reproduce EVM-QuestBench runs and the experimental settings that affect run to run variance. For strict reproducibility, record the fork block height, model sampling parameters, and the task parameter random seed.
A.1 Environment and Dependencies
Experiments require a local EVM mainnet fork and a TypeScript runner.
  • Python 3.10 or newer
  • Node.js 18 or newer (Bun 1.0 or newer is also supported)
  • Foundry with Anvil available in PATH
A.2 Execution Commands
The following commands run the benchmark for a given model identifier.
A.3 Key Experimental Settings
Table 3 lists the settings that most often explain score variance across reruns.
Item Value
Chain EVM-compatible mainnet fork
Runtime TypeScript runner using ethers.js v6
Atomic output One module exporting executeSkill that returns one TransactionRequest
Composite output Iterative calls; each round returns either one tx module or a control JSON (query, error, submit)
Composite rounds Bounded by task optimal_steps and a multiplier such as max_rounds_multiplier
Non determinism Fork block height, model sampling parameters, random parameter sampling when seed is not fixed
Table 3: Reproducibility relevant settings.
Appendix B Prompt and Interface Specification
This section documents the runner interface and response validity rules used by the evaluator. The goal is to make the execution contract between model and runner explicit.
B.1 Runner Interface
Each model outputs a TypeScript module exporting an entry function. The runner provides providerUrl, the agent EOA address, and a contract address map for the local fork.
export async function executeSkill(
providerUrl: string,
agentAddress: string,
deployedContracts: Record<string, string>
): Promise<Record<string, unknown>> {
const tx: Record<string, unknown> = {
to: "0x...",
data: "0x..."
return tx;
B.2 Atomic and Composite Prompt Roles (Schematic)
Atomic tasks require a single transaction that satisfies post execution checks. Composite tasks allow multi round interaction and apply step efficiency decay based on the number of rounds consumed.
Atomic role (schematic)
Produce TypeScript in a code block.
Return exactly one transaction request object.
Use task provided parameters and addresses.
Composite role (schematic)
You may query chain state before executing.
Each round returns either one tx module or a JSON control message.
Completion is signaled with {"submit": true}.
Fewer rounds yield higher score via step efficiency decay.
B.3 Schema Invalid Rules
A response is marked schema_invalid if it cannot be executed under the runner contract.
  1. 1.
    Missing exported executeSkill
  2. 2.
    Function signature mismatch
  3. 3.
    Return value is not a transaction like object
  4. 4.
    Missing required to field
  5. 5.
    Serialization failure under ethers.js
  6. 6.
    No valid TypeScript code block when code is required
  7. 7.
    Control JSON is not parseable in composite control rounds
Appendix C Task Definition Schema
This section summarizes task fields that are most relevant for reproduction and error diagnosis.
C.1 Atomic Task Fields
Atomic tasks specify one on chain action and are validated by post execution constraints.
Field Type Description
id string Unique task identifier
category, subcategory string Task family tags for coverage analysis
difficulty string easy, easy-medium, medium, hard
natural_language_
templates
string[] Instruction templates for prompts
parameters object Typed params with sampling ranges
validation object Post-execution checks and weights
Table 4: Atomic task schema summary.
C.2 Composite Task Fields
Composite tasks add a workflow template and a scoring strategy that emphasizes end to end completion.
Field Type Description
composite_structure object Workflow motif and step sequence
atomic_operations array Ordered atomic steps by atomic_id
optimal_steps int $K_{\text{opt}}$ for step efficiency decay
max_rounds_multiplier int Upper bound on interaction rounds
scoring_strategy object Validator config for end state
Table 5: Composite task schema additions.
Appendix D Validators and Scoring
EVM-QuestBench uses validator based post execution scoring rather than reference code matching. Validators check receipts and post state signals such as balance deltas, allowances, and protocol specific outcomes.
D.1 Validator Families
Table 6 summarizes common validator families and the constraints they evaluate.
Family Typical tolerance Examples of checks
Native and ERC20 transfer small relative tolerance receipt success, recipient, amount, balance delta
Approval exact match spender address, allowance target
DEX swap slippage tolerance router, path, amountIn and minOut, output delta
Liquidity small relative tolerance token amounts, LP minted, pool state updates
Staking and farming small relative tolerance pool identifier, staked balance delta, reward signals
Queries numeric tolerance returned value accuracy, output format correctness
Table 6: Validator families summary.
D.2 Composite Step Efficiency Decay
Composite tasks apply an outcome based base score and multiply it by a step efficiency factor.
$\text{Score}_{\text{final}}=\text{Score}_{\text{base}}\times\min\left(1.0,\frac{K_{\text{opt}}}{K_{\text{act}}}\right).$ (1)
Here $K_{\text{opt}}$ is the task defined optimal step count and $K_{\text{act}}$ is the number of executed rounds, including retries and query rounds.
$K_{\text{opt}}$ $K_{\text{act}}$ Decay Base score 100
3 2 1.00 100
3 3 1.00 100
3 4 0.75 75
3 6 0.50 50
Table 7: Step efficiency decay examples.
D.3 Validator Tolerance Sensitivity
To assess whether the tolerance thresholds influence evaluation outcomes, we analyzed the deviation distribution across 5 rounds $\times$ 20 models $\times$ 107 tasks (10,700 question instances). Table 8 reports the results.
Outcome Count %
Score >0>0, deviation $0\%$ (exact match) 7,013 65.5%
Score =0=0, execution failure / schema error 1,931 18.0%
Score =0=0, transaction revert 493 4.6%
Score =0=0, other failure 1,191 11.1%
Score >0>0, deviation $0\%$ (float precision) 72 0.7%
Table 8: Actual deviation distribution across all 10,700 evaluated question instances (20 models $\times$ 5 rounds $\times$ 107 tasks). Non-zero deviations are all ${\sim}10^{-11}\%$ (floating-point precision), approximately $10^{7}\times$ smaller than the strictest tolerance.
We varied the tolerance multiplier from $0.5\times$ to $3\times$ and re-scored all instances. Under all multiplier settings, all 20 models produce identical scores and rankings (Spearman’s $\rho=1.000$, $\Delta=0\%$). This confirms that tolerances function as safety margins and do not influence reported scores or rankings. The bimodal nature of model outputs — either exactly correct or catastrophically wrong (transaction revert, schema error) — means no question instances fall in the “close but off by a few percent” zone where tolerance thresholds would matter.
Appendix E Additional Analysis Figures
This section provides supporting plots used to interpret difficulty and failure modes.
E.1 Total Score Ranking
Figure 6 shows the total score ranking across all 20 models (avg@5).
Figure 6: Model ranking by total score (avg@5, 20 models).
E.2 Composite Pattern Difficulty
Figure 7 reports pass rates for frequent composite workflow patterns. Lower pass rates indicate higher coordination burden across steps, including prerequisite approvals and parameter propagation.
Figure 7: Pass rate by composite workflow pattern for the ten most frequent patterns.
E.3 Step Overhead Distribution
Figure 8 shows the distribution of step overhead $\Delta=K_{\text{act}}-K_{\text{opt}}$ aggregated across composite runs. Positive overhead indicates extra rounds consumed by redundant queries, retries, or incorrect intermediate actions.
Figure 8: Distribution of step overhead $\Delta=K_{\text{act}}-K_{\text{opt}}$.
E.4 Atomic Subcategory Difficulty
Figure 9 summarizes atomic pass rates by subcategory. This view localizes which single transaction primitives are brittle, such as query heavy tasks or protocol specific calls with strict calldata requirements.
Figure 9: Pass rate by atomic subcategory.
Appendix F Full Evaluation Tables
This section lists full model coverage and step overhead statistics used in the main analysis.
F.1 Quadrant Summary
Table 9 summarizes the four capability profiles based on AtomicAvg and CompositeAvg medians (50.5 and 67.5 respectively).
Profile Count Interpretation
High-High 7 strong precision and workflows
High-Low 3 strong precision, weaker workflows
Low-High 3 weaker precision, stronger workflows
Low-Low 3 weaker on both splits
Atomic-Only 4 composite split near zero
Table 9: Quadrant summary based on AtomicAvg and CompositeAvg medians.
F.2 Model Coverage and Quadrant Classification
Table 10 reports coverage, normalized averages (avg@5), total score, and the quadrant label based on AtomicAvg and CompositeAvg medians.
Model Atomic Composite AtomicAvg CompAvg Total Quadrant
Claude-Sonnet-4.5 62/62 45/45 69.0 88.0 8235.8 High-High
Gemini-3-Pro 62/62 45/45 69.4 79.1 7863.0 High-High
GPT-5 62/62 45/45 66.6 81.0 7773.7 High-High
GPT-5.1 62/62 45/45 59.9 80.4 7331.7 High-High
Kimi-K2-Thinking 62/62 45/45 58.1 80.7 7238.3 High-High
Gemini-2.5-Flash 62/62 45/45 52.7 83.7 7033.4 High-High
GPT-OSS-120B 62/62 45/45 55.8 77.1 6925.7 High-High
Qwen3-235B 62/62 45/45 58.2 47.0 5724.5 High-Low
Claude-Haiku-4.5 62/62 45/45 52.0 52.6 5588.4 High-Low
Mimo-V2-Flash 62/62 45/45 51.4 38.3 4912.3 High-Low
DeepSeek-V3.2 62/62 45/45 49.5 80.9 6709.2 Low-High
Grok-4-Fast 62/62 45/45 48.1 71.3 6187.0 Low-High
GPT-5.1-Mini 62/62 45/45 40.5 70.9 5700.0 Low-High
Qwen3-Max 62/62 45/45 44.8 64.1 5658.5 Low-Low
GLM-4.6 62/62 45/45 46.5 43.8 4852.0 Low-Low
Qwen3-30B 62/62 45/45 25.1 44.7 3565.5 Low-Low
Devstral-2512 62/62 45/45 46.7 0.5 2913.5 Atomic-Only
Qwen3-Coder 62/62 45/45 29.0 3.6 1961.0 Atomic-Only
Qwen3-Coder-Flash 62/62 45/45 3.6 0.8 260.5 Atomic-Only
Qwen3-Coder-30B 62/62 45/45 2.3 0.1 145.0 Atomic-Only
Table 10: Model coverage and quadrant classification (avg@5). Averages are normalized to the 0–100 scale. Medians: AtomicAvg = 50.5, CompositeAvg = 67.5.
F.3 Step Overhead and Decay Impact Summary
Table 11 reports overhead statistics (avg@5) for models with non-trivial composite scores. The overhead statistic is $\Delta=K_{\text{act}}-K_{\text{opt}}$.
Model Mean $\Delta$ Over-opt (%) P90 $\Delta$ Comp. score
DeepSeek-V3.2 +0.05 44 3 3638.4
Kimi-K2-Thinking +0.17 42 3 3633.3
GPT-5 +0.19 48 3 3645.7
Claude-Sonnet-4.5 +0.34 36 2 3958.0
GPT-OSS-120B +0.41 47 3 3468.1
Gemini-2.5-Flash +0.67 45 3 3768.0
GPT-5.1 +0.70 56 3 3617.1
Gemini-3-Pro +0.83 59 3 3561.0
Grok-4-Fast +1.02 62 3 3206.8
GPT-5.1-Mini +1.24 67 3 3189.8
Qwen3-Max +1.35 63 3 2883.5
Qwen3-Coder +2.36 86 4 160.0
Qwen3-235B +2.47 84 5 2115.9
Claude-Haiku-4.5 +2.51 88 5 2367.4
GLM-4.6 +2.55 89 5 1971.0
Mimo-V2-Flash +2.76 92 5 1723.1
Qwen3-30B +2.88 92 5 2011.5
Table 11: Step overhead and decay impact summary (avg@5). Models with near-zero composite scores (Devstral-2512, Qwen3-Coder-Flash, Qwen3-Coder-30B) are omitted.
Appendix G End-to-End Walkthrough
We illustrate the complete evaluation pipeline with one atomic task: bnb_transfer_percentage.
Step 1: Parameter Instantiation.
The runner selects a natural language template and samples numeric parameters:
  • Template selected: “Transfer {percentage}% of my BNB balance to {recipient}.”
  • Parameters sampled: percentage =15=15, recipient == 0xA1b2...C3d4 (randomly generated).
  • Final instruction: “Transfer 15% of my BNB balance to 0xA1b2…C3d4.”
Step 2: Script Generation.
The model receives the instruction along with system context (chain ID, available libraries, account address, contract ABIs). It generates a TypeScript module that:
  1. 1.
    Queries the current BNB balance via provider.getBalance(account).
  2. 2.
    Computes amount = balance * 15n / 100n.
  3. 3.
    Returns a transaction request: {to: recipient, value: amount}.
Step 3: Execution.
The runner signs and submits the transaction on the forked chain. The transaction executes successfully (receipt status =1=1). The runner records pre-execution and post-execution balances for both the sender and recipient.
Step 4: Validation.
The bnb_transfer_percentage validator performs four checks:
  1. 1.
    Transaction Success (30 pts): receipt status =1=1. ✓
  2. 2.
    Recipient Correct (20 pts): tx.to matches the sampled recipient. ✓
  3. 3.
    Transfer Amount (20 pts): actual transferred amount within 0.1% of expected ($\text{balance}\times 15\%$). ✓
  4. 4.
    Balance Change (30 pts): sender balance decreased by amount ++ gas; recipient balance increased by amount, both within 1% tolerance. ✓
Final score: $30+20+20+30=100$ out of 100.
Key design points.
The expected transfer amount is computed dynamically from the fork state (not hardcoded), so the ground truth is always consistent with the execution environment. The tolerance on balance change (1%) absorbs gas cost estimation differences without admitting incorrect transfers.
Appendix H Case Studies
This section provides representative workflows that illustrate common failure modes. The goal is to highlight failure points such as parameter propagation, prerequisite handling, and protocol sequencing.
H.1 Case 1: Complex Multi step DeFi Workflow
Task: composite_complete_swap_stake_workflow.
Workflow: query, approve, swap, add liquidity, stake, verify.
Optimal steps: 6.  Pass rate: 59%.
This task combines prerequisite checks with multiple protocol interactions. Failures often arise from incorrect swap path or slippage, mismatched token ratios during liquidity provision, or staking to an incorrect pool. Models that succeed typically keep parameters consistent across all steps and avoid redundant query rounds that increase $K_{\text{act}}$.
H.2 Case 2: Batch Approval Workflow
Task: composite_batch_approve_2_tokens.
Workflow: approve token A, approve token B, execute the downstream action.
Optimal steps: 3.  Pass rate: 56%.
This pattern stresses multi-transaction sequencing and nonce handling. Common failures include nonce conflicts under rapid submissions, gas estimation errors for back-to-back approvals, and incomplete approvals that allow only one token to proceed.
H.3 Case 3: Natural Language Misinterpretation
Task: bnb_transfer_percentage.
Instruction: “Transfer 15% of my BNB balance to 0xA1b2…C3d4.”
Failure mode: The model interprets “15%” as a fixed amount of 15 BNB rather than 15% of the current balance. This produces a correct transaction that executes successfully but transfers the wrong amount — the validator detects a deviation exceeding the 0.1% tolerance on the Transfer Amount check.
Affected models: Observed in lower-tier models that lack numeric reasoning about percentage-based instructions.
H.4 Case 4: Planning-Level Failure
Task: composite_approve_and_swap.
Instruction: “Swap 0.05 BNB for USDT on PancakeSwap.”
[…]
Impact: The model scores 0 on this task. Step-efficiency decay is irrelevant since the base score is 0.
H.5 Case 5: Partial Execution in Multi-Step Workflows
Task: composite_complete_swap_stake_workflow.
Instruction: “Swap BNB for CAKE, add CAKE-BNB liquidity, then stake the LP tokens.”
Failure mode: The model successfully completes the swap and liquidity addition but fails on the staking step due to an incorrect staking pool address or mismatched LP token. The validator checks the final state (LP tokens staked) and assigns a base score of 0 despite partial progress.
Observation: This pattern accounts for the majority of composite failures among mid-tier models, where individual actions succeed but end-to-end workflow orchestration breaks down.